OpenStudy (anonymous):

Q one- >10^-12=? (no clue for this one) Q two ->3^5+4^12=? would it be 3.3.3.3.3+4.4.4.4.4.4.4.4.4.4.4.4= 16,777,459 or 7^17 or either lol

OpenStudy (anonymous):

For the example they show that is like 10^-12....(Example is 10^-4) they do 1/10.10.10.10 and turn that answer into a decimal... and Question two is 3^5+ 4^12=???

OpenStudy (anonymous):

My error: 10^-12 is 1/(10*10*...*10) = (0.1)^12 = (1/10)^12 Your approach to Q2 is correct: 3*3*3*3*3 + etc.

OpenStudy (anonymous):

So Q2 would be 3^5+4^12= 16,777,459 or 7^17? I haven't done this stuff in years trying to help my little cousin... lol